[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Aug 10, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Aug 10 19:24:31 UTC 2017


HSU Shift Log for Thu Aug 10 2017 15:24:24 GMT-0400 (EDT)

Shift: H

Forecaster: R Pasch

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues: Pablo conducted more National TCV training with Dave Z  
and Dave R.

Other Techincal issues: Hot line working very poorly and making lots of  
noise; submitted NCO help desk ticket.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170810/dce896b3/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list